# On the CP Nature of the ‘95 GeV’ Anomalies

### v2

@@ Abstract
Under the assumption that the various evidences of a ‘95 GeV’ excess, seen in data at the Large Electron Positron (LEP) collider as well as the Large Hadron Collider (LHC), correspond to actual signals of new physics Beyond the Standard Model (BSM), we characterise the underlying particle explaining these in terms of its Charge/Parity (CP) quantum numbers in a model independent way. In doing so, we assume the new object having spin-0 and test the CP-even (scalar) and CP-odd (pseudoscalar) hypotheses as well superpositions of these in its $\tau^{+}\tau^{-}$ decays. We prove that the High Luminosity LHC (HL-LHC) will be in a position to disentangle the CP nature of such a new particle.
@@ 1 Introduction
Since the discovery of the 125 GeV Higgs boson at the Large Hadron Collider (LHC) in 2012 ATLAS:2012yve ; CMS:2012qbp , significant efforts have been made by the ATLAS and CMS collaborations in ascertaining the nature of this particle, that is, measuring its width, couplings, spin and CP quantum numbers. All corresponding measurements are pointing towards the discovered Higgs boson being the Standard Model (SM) one. Despite all this, though, the search for physics Beyond the SM (BSM) continues, spurred, on the theoretical side by its many flaws and, on the experimental side, by the ever increasing precision of the analyses as well as additional data sets being collected during Run 3 of the LHC.
@@
The 2-Higgs Doublet Model (2HDM) Gunion:1992hs ; Branco:2011iw , as embedded in various BSM scenarios, including fundamental ones like Supersymmetry Moretti:2019ulc and Compositeness DeCurtis:2018zvh ; DeCurtis:2018iqd , is particularly interesting, as it offers in its particle spectrum both a CP-even and a CP-odd (neutral) Higgs boson, alongside the discovered SM-like Higgs state, both of which can in fact be lighter or heavier than the latter [^1] There are also charged Higgs bosons.. Hence, this BSM scenario has been studied extensively since its inception. However, its general version allows for non-diagonal Yukawa couplings, potentially leading to large Flavor Changing Neutral Currents (FCNCs) at tree level, contrary to experimental evidence. To constrain the latter, from the theoretical side, a $Z_{2}$ symmetry (which can be softly broken, in fact) is typically imposed to define the coupling structure of the two Higgs doublets to SM fermions. This classification includes the so-called Type-I, Type-II, lepton-specific and flipped scenarios Branco:2011iw , alongside the 2HDM Type-III, which allows direct couplings of both doublets to all SM fermions. The 2HDM Yukawa structure is then refined by both theoretical consistency requirements and experimental measurements of the $h$ mass and couplings or limits on the the same parameters for the $H,A$ and $H^{\pm}$ states.
@@
Given that a 2HDM, as mentioned, allows (in some Types) for neutral Higgs bosons states ligther than 125 GeV, both ATLAS and CMS have pursued this possibility through dedicated searches in presence of a mass configuration wherein $m_{h}$ and/or $m_{A}<m_{H}=125$ GeV (with $m_{H^{\pm}}$ also rather light, to comply with various high-precision data). (Hereafter, in our notation, the CP-even $H$ state is the SM-like one already discovered while the $h$ and $A$ states are the (light) CP-even and CP-odd Higgs bosons, respectively, plus $H^{\pm}$ are the charged ones.)
@@
Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Bhattacharya:2023lmu ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn ; Arhrib:2024wjj ; Gao:2024qag have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a non-standard Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC. In the attempt to explain the excesses in the $\gamma\gamma$ and $b\overline{b}$ channels, it was found in Refs. Benbrik:2022azi ; Benbrik:2022tlg ; Belyaev:2023xnv ; Belyaev:2024lah ; Benbrik:2024ptw that the 2HDM Type-III with a particular Yukawa texture can successfully accommodate both measurements simultaneously with the lightest CP-even Higgs boson of the model, while being consistent with all relevant theoretical and experimental constraints. In fact, also a 2HDM Type-I can afford similar explanations to the excesses Khanna:2024bah . Further recent studies have shown that actually all three aforementioned signatures can be simultaneously explained in the 2HDM plus a real (N2HDM) Biekotter:2022jyr or a complex (S2HDM) Biekotter:2023jld ; Biekotter:2023oen singlet.
@@
The purpose of this paper is to assume that, indeed, a spin-0 particle of yet undefined CP is behind all such anomalies. However, amongst the latter, we focus here on the $\tau^{+}\tau^{-}$ one, as this has some diagnostic power in terms of pinning down its CP properties, as shown in Refs. Bower:2002zx ; Desch:2003rw ; Berge:2013jra ; Berge:2008wi ; Berge:2008dr ; Berge:2011ij ; Berge:2014sra ; Berge:2014wta ; Berge:2015nua ; Han:2016bvf ; Antusch:2020ngh ; Dutta:2021del ; Esmail:2024jdg . So, let us spend some time on this particular anomaly. The CMS collaboration has reported such an excess in the $\tau^{+}\tau^{-}$ final state across all $\tau$ decay modes (leptonic and hadronic) for both leptons. The local and global significances of this particular excess are $2.6\sigma$ and $2.3\sigma$ for $m_{\tau^{+}\tau^{-}}=95$ GeV, respectively. At $m_{\tau^{+}\tau^{-}}=100$ GeV, these values are $3.1\sigma$ and $2.7\sigma$. By introducing an additional single neutral narrow resonance $h$, the best fit values to the signal rate for this excess are CMS-PAS-HIG-21-001 :
$\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})=7.7^{+3.9}_{-3.1}~{}\text{pb}\quad\text{for}~{}m_{h}=95~{}\text{GeV},$ (1)
$\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})=5.8^{+2.4}_{-2.0}~{}\text{pb}\quad\text{for}~{}m_{h}=100~{}\text{GeV}.$ (2)
In our work, we assume a 95 GeV resonance. Interpetting Eq. (1) in terms of a signal strength, we get
$\mu^{\text{exp}}_{\tau^{+}\tau^{-}}=\frac{\sigma^{\text{exp}}(gg\to h\to\tau^{+}\tau^{-})}{\sigma^{\text{SM}}(gg\to H\to\tau^{+}\tau^{-})}=1.2\pm 0.5,$ (3)
@@
where we use a symmetric uncertainty interval for such a signal strength, which is obtained from the lower uncertainty interval of the quoted result for $\sigma(gg\to h)\;{\rm BR}(h\to\tau^{+}\tau^{-})$. Also, in the signal strength, $H$ represents a fictitious SM Higgs boson with mass 95 GeV.
@@
$\mu^{\text{exp}}_{\gamma\gamma}=\frac{\sigma^{\text{exp}}(gg\to h\to\gamma\gamma)}{\sigma^{\text{SM}}(gg\to H\to\gamma\gamma)}=0.6\pm 0.2$ (4)
@@
$\mu^{\text{exp}}_{bb}=\frac{\sigma^{\text{exp}}(e^{+}e^{-}\to ZS\to Zb\bar{b})}{\sigma^{\text{SM}}(e^{+}e^{-}\to Zh\to Zb\bar{b})}=0.117\pm 0.057$ (5)
@@
In this work, in order to ascertain the CP state of the underlying spin-0 resonance (which, for illustrative purposes, we also assume to be a Higgs state), we have adopted a model agnostic explanation to all the excesses. As we particularly focus on the CP nature of this 95 GeV object, we have setup a field with Yukawa couplings to all fermions, i.e., proportional to their mass, scaled by the corresponding SM Yukawa, while allowing in such an interaction for the simultaneous presence of both a $\gamma_{5}$ (i.e., a pseudoscalar interaction) and a unit matrix (i.e., a scalar interaction) structure, in both the $ht\bar{t}$ and $h\tau^{+}\tau^{-}$ vertices, entering in production and decay, respectively. After exploring the parameter space of this simplified model by testing it against theoretical and experimental constraints, including fitting the described excesses, we look at unravelling the CP properties of the $h$ state at the LHC in the $\tau^{+}\tau^{-}$ channel using as diagnostic observable the signed acoplanarity angle $\phi_{\rm CP}$ between the two $\tau$ lepton decay planes. In order to reconstruct the latter, we have used one-prong hadronic $\tau$ decays and Impact Parameter (IP) information Berge:2014sra ; Berge:2015nua , wherein for the former we limited ourselves to the case of $\pi$ and $\rho$ mesons. By then performing a robust detector level simulation, we show that the it is indeed possible to unambiguously differentiate a CP-even from a CP-odd as well as a maximally CP-mixed signal at the LHC with 3000 fb-1 of data per experiment, which is the expected final dataset size of the High-Luminosity LHC (HL-LHC) Gianotti:2002xx ; Apollinari:2015wtw . However, we deem this result a conservative one since, as we gather more data at the current LHC (Run 3), the significance of the aforementioned excesses may well grow further, thereby stimulating dedicated analyses aimed at accessing the CP properties of this new resonance which may well prove to offer a better sensitivity than the one worked out in this article.
@@ 2 A Simplified Model to Address the Excesses
We are interested in the study of CP properties of the observed spin-0 resonance, which might come from various UV complete scenarios like the Two Higgs Doublet Model, Next-to Minimal Supersymmetric scenarios (NMSSM) and other scalar extensions of the Standard Model. To remain model agnostic, we assumed a simplified model where, in addition to the SM spectrum, we consider only an additional scalar ($S$) of mass 95 GeV. The scalar particle $S$ couples with the fermions as follows:
@@
We have introduced CP-Violating (CPV) Yukawa coupling with $\alpha$ being the CPV parameter. $\alpha$ will be zero ($\pi/2$) when the scalar $S$ is a CP-even (CP-odd) particle. For simplicity, we have assumed a uniform $CP$ mixing among generations. The dimensionless real parameters $\rho^{S}_{f}$ ($f=t,b,\tau$) are the Yukawa coupling modifiers. The observed SM-like $CP$-even Higgs boson ($h$) of mass 125 GeV can couple with $S$ depending on the UV completion. Our analysis will not depend on the coupling of $S$ with the SM Higgs boson.
The coupling of the scalar $S$ with the weak neutral gauge boson, denoted as $g_{SVV}$, does not affect the LHC excess observed in $\tau^{+}\tau^{-}$ and $\gamma\gamma$ channel. However, a non-zero $g_{SVV}$ is necessary to explain the LEP $b\bar{b}$ excess. Such coupling can appear if $S$ is part of a weak multiplet which mixes with the SM Higgs doublet. Since the observed 125 GeV Higgs boson at the LHC closely resembles the SM, the mixing of the Higgs doublet can not be too large, and we have considered $g_{SVV}<0.3$.
@@ 2.1 Production and Decay Modes
Considering the large top mass and presuming $\rho^{h}_{b}$ not to be significantly large compared to $\rho^{h}_{t}$, we can obtain the gluon effective coupling described by a dimension five operator
@@
where $G_{\mu\nu}$ is the gluon field strength tensor and $\widetilde{G}_{\mu\nu}=\frac{1}{4}\epsilon_{\mu\nu\rho\sigma}G^{\rho\sigma}$ is the corresponding dual field strength tensor and $A_{S}$ and $A^{\prime}_{S}$ are effective couplings. We incorporated Eq. (6) in FeynRules Christensen:2008py ; Alloul:2013bka and used SusHi Harlander:2012pb ; Harlander:2016hcx in order to estimate the NNLO production cross-section via gluon fusion. The gluon fusion cross section at the large top mass limit at NNLO is given by
$\sigma(pp\to gg\to S)={\rho_{t}^{S}}^{2}\Big{(}76.35\cos^{2}\alpha+176.32\sin^{2}\alpha\Big{)}~{}\text{pb},$ (8)
where the first term is the contribution from the $CP$ even part, and the second is the $CP$ odd contribution. The numerical coefficients are obtained from SusHi Harlander:2012pb ; Harlander:2016hcx .
For our analysis we have considered that the scalar $S$ can decay via the following channels Choi:2021nql ,

### v3

@@ Abstract
Under the assumption that the various evidences of a ‘95 GeV’ excess, seen in data at the Large Electron Positron (LEP) collider as well as the Large Hadron Collider (LHC), correspond to actual signals of new physics Beyond the Standard Model (BSM), we characterise the underlying particle explaining these anomalies in terms of its Charge/Parity (CP) quantum numbers. In doing so, we use $\chi^{2}$ fits to test the CP-even (scalar) and CP-odd (pseudoscalar) hypotheses and superpositions of these, thus under the assumption of a spin-0 resonance. This is done through the exploitation of $\tau^{+}\tau^{-}$ decays, in both their fully hadronic and semi-leptonic modes, in a model-independent way, so that our approach enables one to test a variety of BSM hypotheses, having proven here that the High-Luminosity LHC (HL-LHC) will be in a position to disentangle the CP nature of such a new particle within $\pm(0.27-0.47)$ radians of the true hypothesis at $90\%$ Confidence Level (CL), depending on the assumed background systematics.
preprint: CQUeST-2025-0763
@@ 1 Introduction
Since the discovery of the 125 GeV Higgs boson at the Large Hadron Collider (LHC) in 2012 ATLAS:2012yve ; CMS:2012qbp , significant efforts have been made by the ATLAS and CMS collaborations in ascertaining the nature of this particle, that is, measuring its width, couplings, spin and CP quantum numbers. All corresponding measurements are pointing towards the discovered Higgs boson being the Standard Model (SM) one. Despite all this, though, the search for physics Beyond the SM (BSM) continues, spurred, on the theoretical side, by its many flaws and, on the experimental side, by the ever increasing precision of the analyses as well as additional data sets being collected during Run 3 of the LHC.
@@
The 2-Higgs Doublet Model (2HDM) Gunion:1992hs ; Branco:2011iw , as embedded in various BSM scenarios, including fundamental ones like Supersymmetry Moretti:2019ulc and Compositeness DeCurtis:2018zvh ; DeCurtis:2018iqd , is particularly interesting, as it offers in its particle spectrum both a CP-even and a CP-odd (neutral) Higgs boson, alongside the discovered SM-like Higgs state, both of which can in fact be lighter or heavier than the latter [^1] The CP-even (neutral) Higgs states of the 2HDM are labelled as $h$ and $H$ (with $m_{h}<m_{H}$) while the CP-odd (neutral) one as $A$, plus there are also two charged Higgs bosons, $H^{\pm}$.. Hence, this BSM scenario has been studied extensively since its inception. However, its general version allows for non-diagonal Yukawa couplings, potentially leading to large Flavor Changing Neutral Currents (FCNCs) at tree level, contrary to experimental evidence. To constrain the latter, from the theoretical side, a $Z_{2}$ symmetry (which can be softly broken, in fact) is typically imposed to define the coupling structure of the two Higgs doublets to SM fermions. This classification includes the so-called Type-I, Type-II, lepton-specific and flipped scenarios Branco:2011iw , alongside the 2HDM Type-III, which allows direct couplings of both doublets to all SM fermions. The 2HDM Yukawa structure is then refined by both theoretical consistency requirements and experimental measurements of the SM-like Higgs mass and couplings or limits on the same parameters for the other Higgs states.
@@
Given that a 2HDM, as mentioned, allows (in some Types) for neutral Higgs bosons states lighter than 125 GeV, both ATLAS and CMS have pursued this possibility through dedicated searches in presence of a mass configuration wherein $m_{h}$ and/or $m_{A}<m_{H}=125$ GeV (with $m_{H^{\pm}}$ also rather small, to comply with various high-precision data). (Hereafter, in our notation, when referring to the 2HDM, the CP-even $H$ state is the SM-like one already discovered while the $h$ and $A$ states are the lightest CP-even and CP-odd Higgs bosons, respectively.) However, in performing such searches, neither ATLAS nor CMS have defined potential signal regions based on 2HDM assumptions, as therein the latter only entered at the interpretation level. Therefore, their results can be used in a model-independent way, which is what we will be doing here by introducing a simplified model approach wherein the underlying BSM scenario is parameterized solely in terms of masses and (gauge plus Yukawa) couplings that are fitted to data, yet, we assume the same particle content of the generic 2HDM when computing the decay width of a new Higgs state that we will introduce (hereafter, denoted by $S$).
@@
Since all the excesses appear in very similar mass regions, several studies Cao:2016uwt ; Heinemeyer:2021msz ; Biekotter:2021qbc ; Biekotter:2019kde ; Cao:2019ofo ; Biekotter:2022abc ; Iguro:2022dok ; Li:2022etb ; Cline:2019okt ; Biekotter:2021ovi ; Crivellin:2017upt ; Cacciapaglia:2016tlr ; Abdelalim:2020xfk ; Biekotter:2022jyr ; Biekotter:2023jld ; Azevedo:2023zkg ; Biekotter:2023oen ; Cao:2024axg ; Wang:2024bkg ; Li:2023kbf ; Dev:2023kzu ; Borah:2023hqw ; Cao:2023gkc ; Ellwanger:2023zjc ; Aguilar-Saavedra:2023tql ; Ashanujjaman:2023etj ; Bhattacharya:2023lmu ; Dutta:2023cig ; Ellwanger:2024txc ; Diaz:2024yfu ; Ellwanger:2024vvs ; Ayazi:2024fmn ; Arhrib:2024wjj ; Gao:2024qag have explored the possibility of simultaneously explaining these anomalies within BSM frameworks featuring a Higgs state lighter than 125 GeV, while being in agreement with current measurements of the properties of the SM-like Higgs state observed at the LHC. In the attempt to explain the excesses in the $\gamma\gamma$ and $b\overline{b}$ channels, it was found in Refs. Benbrik:2022azi ; Benbrik:2022tlg ; Belyaev:2023xnv ; Belyaev:2024lah ; Benbrik:2024ptw that the 2HDM Type-III with a particular Yukawa texture can successfully accommodate both measurements simultaneously with the lightest CP-even Higgs boson of the model, while being consistent with all relevant theoretical and experimental constraints. In fact, also a 2HDM Type-I can afford similar explanations to the excesses Khanna:2024bah . Further recent studies have shown that actually all three aforementioned signatures can be simultaneously explained in the 2HDM plus a real (N2HDM) Biekotter:2022jyr or a complex (S2HDM) Biekotter:2023jld ; Biekotter:2023oen singlet.
@@
The purpose of this paper is to assume that, indeed, a spin-0 particle of yet undefined CP is behind all such anomalies. However, amongst the latter, we focus here on the $\tau^{+}\tau^{-}$ one, as this has some diagnostic power in terms of pinning down its CP properties, as shown in Refs. Bower:2002zx ; Desch:2003rw ; Berge:2013jra ; Berge:2008wi ; Berge:2008dr ; Berge:2011ij ; Berge:2014sra ; Berge:2014wta ; Berge:2015nua ; Han:2016bvf ; Antusch:2020ngh ; Dutta:2021del ; Esmail:2024jdg . So, let us spend some time on this particular anomaly. The CMS collaboration has reported such an excess in the $\tau^{+}\tau^{-}$ final state across all $\tau$ decay modes (leptonic and hadronic) for both leptons. The local and global significances of this particular excess are $2.6\sigma$ and $2.3\sigma$ for $m_{\tau^{+}\tau^{-}}=95$ GeV, respectively. At $m_{\tau^{+}\tau^{-}}=100$ GeV, these values are $3.1\sigma$ and $2.7\sigma$. By introducing an additional single neutral narrow resonance $S$, the best fit values to the signal rate for this excess are CMS-PAS-HIG-21-001 :
$\sigma(gg\to S)\;{\rm BR}(S\to\tau^{+}\tau^{-})=7.7^{+3.9}_{-3.1}~\text{pb}\quad\text{for}~m_{S}=95~\text{GeV},$ (1)
$\sigma(gg\to S)\;{\rm BR}(S\to\tau^{+}\tau^{-})=5.8^{+2.4}_{-2.0}~\text{pb}\quad\text{for}~m_{S}=100~\text{GeV}.$ (2)
In our work, we assume a 95 GeV resonance. Interpreting Eq. (1) in terms of a signal strength, we get
$\mu^{\text{exp}}_{\tau^{+}\tau^{-}}=\frac{\sigma^{\text{exp}}(gg\to S\to\tau^{+}\tau^{-})}{\sigma^{\text{SM}}(gg\to\Phi\to\tau^{+}\tau^{-})}=1.2\pm 0.5,$ (3)
@@
where we use a symmetric uncertainty interval for such a signal strength, which is obtained from the lower uncertainty interval of the quoted result for $\sigma(gg\to S)\;\times\;{\rm BR}(S\to\tau^{+}\tau^{-})$. Also, in the signal strength, $\Phi$ represents a fictitious SM Higgs boson with mass 95 GeV.
@@
$\mu^{\text{exp}}_{\gamma\gamma}=\frac{\sigma^{\text{exp}}(gg\to S\to\gamma\gamma)}{\sigma^{\text{SM}}(gg\to\Phi\to\gamma\gamma)}=0.6\pm 0.2$ (4)
@@
$\mu^{\text{exp}}_{bb}=\frac{\sigma^{\text{exp}}(e^{+}e^{-}\to ZS\to Zb\bar{b})}{\sigma^{\text{SM}}(e^{+}e^{-}\to Z\Phi\to Zb\bar{b})}=0.117\pm 0.057$ (5)
@@
In this work, in order to ascertain the CP state of the underlying spin-0 resonance (which, for illustrative purposes, we also assume to be a Higgs state), we have adopted a model-agnostic explanation to all the excesses. As we particularly focus on the CP nature of this 95 GeV object, we have setup a field with Yukawa couplings to the third generation fermions, scaled by the corresponding SM Yukawa, while allowing in such an interaction for the simultaneous presence of both a $\gamma_{5}$ (i.e., a pseudoscalar interaction) and a unit matrix (i.e., a scalar interaction) structure, in both the $St\bar{t}$ and $S\tau^{+}\tau^{-}$ vertices, entering in production and decay, respectively. After exploring the parameter space of this simplified model by testing it against theoretical and experimental constraints, including fitting the described excesses, we look at unravelling the CP properties of the $S$ state at the LHC in the $\tau^{+}\tau^{-}$ channel using as diagnostic observable the signed acoplanarity angle $\phi_{\rm CP}$ between the two $\tau$ lepton decay planes. In order to reconstruct the decay planes, we have used both hadronic and semi-leptonic decays of the tau lepton, and used the Impact Parameter (IP) information Berge:2014sra ; Berge:2015nua of the charged prong as well as the $\rho$-decay method Bower:2002zx ; Berge:2015nua ; Hansen:2020ecn . By then performing a robust detector-level simulation, we show that it is indeed possible to unambiguously differentiate a CP-even from a CP-odd as well as a maximally CP-mixed signal at the LHC with 3000 fb-1 of data per experiment, which is the expected final dataset size of the HL-LHC Gianotti:2002xx ; Apollinari:2015wtw . However, we deem this result a conservative one since, as we gather more data at the current LHC (Run 3), the significance of the aforementioned excesses may well grow further, thereby stimulating especially dedicated analyses aimed at accessing the CP properties of this potential new resonance, which may well prove to offer an even better sensitivity than the one established in this article. Yet, our findings are altogether already very encouraging in this respect. For the observed 125 GeV scalar, a similar CP characterization has been done by ATLAS ATLAS:2022akr and CMS collaboration CMS:2021sdq . However, at the lower mass, the signal rates, backgrounds as well as kinematics are different, and no dedicated CP study exists in this mass region. We believe this is important in order to test various hypotheses made in the literature about the CP properties of possible interpretations.
@@ 2 A Simplified Model to Address the Excesses
We are interested in the study of CP properties of the observed spin-0 resonance, which might come from various UV complete scenarios like the 2HDM, Next-to Minimal Supersymmetric Standard Model (NMSSM) and other scalar extensions of the SM. To remain model agnostic, we assumed a simplified model where, in addition to the SM spectrum, we consider only an additional scalar ($S$) of mass 95 GeV. The scalar particle $S$ couples with the third generation fermions as follows:
@@
We have introduced CP-Violating (CPV) Yukawa coupling with $\alpha$ being the CPV parameter. Here, $\alpha$ will be zero ($\pi/2$) when the scalar $S$ is a CP-even (CP-odd) Higgs particle (e.g., in the 2HDM, this will be $H$ (AA)). For simplicity, we have assumed a uniform CP mixing among generations. The dimensionless real parameters $\rho^{S}_{f}$ ($f=t,b,\tau$) are the Yukawa coupling modifiers. The observed SM-like CP-even Higgs boson of mass 125 GeV can couple with $S$ depending on the Ultra-Violet (UV) completion. Our analysis will not depend on the coupling of $S$ with the SM Higgs boson.
The coupling of the scalar $S$ with the weak neutral gauge boson, denoted as $g_{SVV}$, does not affect the LHC excess observed in the $\tau^{+}\tau^{-}$ and $\gamma\gamma$ channels. However, a non-zero $g_{SVV}$ is necessary to explain the LEP $b\bar{b}$ excess. Such coupling can appear if $S$ is part of a weak multiplet which mixes with the SM Higgs doublet. Since the observed 125 GeV Higgs boson at the LHC closely resembles the SM one, the mixing of the Higgs doublet cannot be too large, and we have considered $g_{SVV}<0.3$.
@@ 2.1 Production and Decay Modes
Considering the large top mass and presuming $\rho^{S}_{b}$ not to be significantly large compared to $\rho^{S}_{t}$, we can obtain the gluon effective coupling described by a dimension five operator
@@
where $G_{\mu\nu}$ is the gluon field strength tensor and $\widetilde{G}_{\mu\nu}=\frac{1}{4}\epsilon_{\mu\nu\rho\sigma}G^{\rho\sigma}$ is the corresponding dual field strength tensor and $A_{S}$ and $A^{\prime}_{S}$ are effective couplings. We incorporated Eq. (6) in FeynRules Christensen:2008py ; Alloul:2013bka and used SusHi Harlander:2012pb ; Harlander:2016hcx in order to estimate the Next-to-Next-to-Leading Order (NNLO) production cross section via gluon fusion. The gluon fusion cross section in the large top mass limit at NNLO is given by
$\sigma(gg\to S)={\rho_{t}^{S}}^{2}\Big{(}76.35\cos^{2}\alpha+176.32\sin^{2}\alpha\Big{)}~\text{pb},$ (8)
where the first term is the contribution from the CP-even part, and the second is the CP-odd contribution. The numerical coefficients are obtained from SusHi Harlander:2012pb ; Harlander:2016hcx .
Finally, for our analysis, we have considered that the scalar $S$ can decay via the following channels Choi:2021nql :